OpenStudy (anonymous):

Simplify. 5 square root of 11 end root minus 12 square root of 11 end root minus 2 square root of 11

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (kirbykirby):

\(5\sqrt{11}-12\sqrt{11}-2\sqrt{11}\) ?